OpenStudy (anonymous):

help pweaseeeeeeeeeeeeeeeeee i have the answer i just need to know how to do it The January 14, 2002 issue of TIME magazine reported that Wal-Mart’s 3rd quarter revenues for 2001 were 600% more than its No. 2 competitor, K-Mart. What were K-Mart’s revenues during this period if Wal-Mart’s revenues were $48 billion?

OpenStudy (anonymous):

For the number to be 600% more, WalMart's number was 7 times more than K-Mart's number. 48 = 7 * K K = 48 / 7 = 6.86 billion